assignment 1} add two vectors on a piece of paper or graph paper A] vector 1: 10 Newtons North. B] vector 2: 5 N WSW
1
Expert's answer
2020-04-15T10:32:57-0400

As per the given question,

vector 1 is "\overrightarrow{v_1}=10\hat{j}" is representing as the north

vector 2 is "\overrightarrow{v_2}=-5\cos 45 \hat{i}-5\sin45\hat{j}"

Now, adding two vector,

"\overrightarrow{V}=\overrightarrow{v_1}+\overrightarrow{v_2}=5\cos 45 \hat{i}-5\sin45\hat{j}+10\hat{j}"

"\overrightarrow{V}=\dfrac{-5}{\sqrt{2}}\hat{i}+(10-\dfrac{5}{\sqrt{2}})\hat{j}"